OpenStudy (anonymous):

Solve: 3(x + 4) greater then or equal to x - 2 x greater then or equal to -7 x greater then or equal to 7 x greater then or equal to -7 x greater then or equal to -3

OpenStudy (anonymous):

solve just like you would if it was an equal sign \[3(x+4)\geq x-2\] \[3x+12\geq x-2\] \[2x+12\geq-2\] \[2x\geq-14\] one more step answer above is wrong

OpenStudy (anonymous):

are u sure ? please explain @fernando22

OpenStudy (anonymous):

@fernando22 is wrong, ignore it

OpenStudy (anonymous):

okay, then can u further explain ? i dont know how to do this at all.

OpenStudy (anonymous):

--_--

OpenStudy (anonymous):

divide both sides by 2

OpenStudy (anonymous):

ya tell me .....

OpenStudy (anonymous):

i got x greater then or equal to -7

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yay :) can u help with my other question ? fernando is looking at it and i dont want another wrong answer :/